# Molecular Seeds of Shear: An operator-level necessity result for first-order Chapman–Enskog deviatoric stress

### v1

@@ 6 Main theorem
We now state the principal result.
Theorem 6.1 (Molecular seed of shear).
Assume A1–A8 and the operator bounds of Lemmas 5.15.3. Let $\{f_{\varepsilon}\}_{0<\varepsilon<\varepsilon_{0}}$ be a family of sufficiently regular solutions of the rescaled Boltzmann equation under Chapman–Enskog scaling on the domain specified in A2, and write the Chapman–Enskog expansion
@@
with $f^{(0)}(x,v,t)=M[\rho,u,T]$. Assume the remainder satisfies the uniform bound
$\|R_{\varepsilon}\|_{L^{2}_{x}L^{2}_{v}(M^{-1})}\leq C\varepsilon^{2}$ (6.2)
for $0<\varepsilon<\varepsilon_{0}$ under the norms and regularity in A3–A7. All assertions of Theorem 6.1 are local-in-time on any interval $[0,T]$ on which the macroscopic fields remain in $H^{s}_{x}$ with the uniform bounds of A7. Constants appearing in (6.2)–(6.6) depend explicitly on $\|(\rho,u,T)\|_{H^{s}_{x}}$, the spectral-gap or hypocoercivity constants, the uniform pseudoinverse bound $C_{\mathrm{inv}}$, and the domain geometry. Uniqueness of $f^{(1)}$ is to be understood pointwise in $x$ in the chosen weighted space under the orthogonality conditions (6.5); potential initial-layer corrections that violate (6.5) are treated as part of the parallel remainder and are controlled by the estimates in Section B. Then:
@@
    (Constitutive decomposition) The pressure tensor admits the asymptotic decomposition
@@
    The mapping $f^{(1)}\mapsto\tau^{(1)}$ is linear and bounded under the prescribed velocity-moment weights.
@@
    (Vanishing for spatially uniform Maxwellians) If $f^{(0)}(x,v,t)$ is spatially uniform (so that $(\partial_{t}^{(0)}+v\cdot\nabla_{x})f^{(0)}\equiv 0$ pointwise in $x$) and the solvability/Fredholm conditions of A6 hold, then $f^{(1)}(x,v,t)\equiv 0$ for all $(x,t)$ and consequently $\tau^{(1)}\equiv 0$. Hence, in this closed, unforced setting there is no $O(\varepsilon)$ deviatoric stress emerging from the Chapman–Enskog expansion.
@@
    (Uniqueness and representation) Conversely, whenever the solvability orthogonality
@@
    holds for all collision invariants $\phi(v)$, the first correction is uniquely determined pointwise in $x$ by the bounded pseudoinverse:
@@
    where $L_{x}$ denotes the linearized operator at the local Maxwellian $M[\rho(x,t),u(x,t),T(x,t)]$. The map $f^{(0)}\mapsto f^{(1)}$ is continuous in the weighted norms specified in A3–A7 with operator norm controlled by the inversion constant $C_{\mathrm{inv}}$ from Lemma 5.2.
Corollary 6.2 (Deterministic necessity of kinetic correction for $O(\varepsilon)$ shear).
@@ 7 BGK worked example
For the BGK model
@@
seek $f=M+\tau f^{(1)}+O(\tau^{2})$. The $O(1)$ equation yields $f^{(0)}=M$. At $O(\tau)$ we have
$(\partial_{t}+v\cdot\nabla_{x})M=-f^{(1)}.$ (7.2)
Evaluating the contribution of $f^{(1)}$ to the pressure tensor at a point $x_{0}$, applying a Galilean shift so $u(x_{0})=0$, and using standard Gaussian integrals for a Maxwellian
@@
we use moments
@@
Then
$\tau^{(1)}_{ij}=\int v_{i}v_{j}f^{(1)}\,dv=-\int v_{i}v_{j}(\partial_{t}+v\cdot\nabla_{x})M\,dv,$ (7.6)
and collecting terms yields the viscous constitutive relation
@@
with $S_{ij}=\tfrac{1}{2}(\partial_{i}u_{j}+\partial_{j}u_{i})-\tfrac{1}{d}(\nabla\cdot u)\,\delta_{ij}$. The Gaussian integrals and prefactors are evaluated under the Galilean shift $u(x_{0})=0$. The BGK remainder estimate is valid in the same weighted $L^{2}_{v}(M^{-1})$ space and the constant in the $O(\varepsilon^{2})$ bound depends explicitly on $\tau$ and $\|(\rho,u,T)\|_{H^{s}_{x}}$.
8 Energy budget, linear stability and transient growth
We summarize the connection between kinetic seed amplitude and macroscopic amplification.
8.1 Kinetic-energy balance
Multiplying the macroscopic momentum equation by $u$ and integrating yields (compressible form suppressed)
$\frac{d}{dt}\int\tfrac{1}{2}\rho|u|^{2}\,dx=-\int\tau_{ij}\partial_{i}u_{j}\,dx+W_{\mathrm{ext}}+B_{\mathrm{boundary}}.$ (8.1)
Substituting the constitutive law $\tau=-\mu S(u)$ recovers viscous dissipation.
8.2 Linearization about a shear profile
For incompressible Navier–Stokes with base flow $U(y)\mathbf{e}_{x}$ and perturbations $u^{\prime}$, the linearized equation is
$\partial_{t}u^{\prime}+U(y)\partial_{x}u^{\prime}+v^{\prime}U^{\prime}(y)\mathbf{e}_{x}=-\nabla p^{\prime}+\nu\Delta u^{\prime},$ (8.2)
with perturbation energy equation
$\frac{d}{dt}E(t)=-\int u^{\prime}_{x}u^{\prime}_{y}U^{\prime}(y)\,dx\,dy\,dz-\nu\int|\nabla u^{\prime}|^{2}\,dV,\qquad E(t)=\tfrac{1}{2}\int|u^{\prime}|^{2}\,dV.$ (8.3)
The production term admits transient amplification for suitable phase relations between $u^{\prime}_{x}$ and $u^{\prime}_{y}$.
8.3 Non-normality and transient growth
The linear operator is non-self-adjoint; the operator norm $\|e^{Lt}\|$ can be large for finite $t$ even if spectral abscissa is negative. Define
$G_{\max}=\sup_{t\geq 0}\|e^{Lt}\|_{L^{2}\to L^{2}}^{2},$ (8.4)
and note $G_{\max}$ can scale like $\mathrm{Re}^{2}$ for some flows [6, 7].
8.4 Linking seed amplitude to threshold
Let $a_{0}$ denote amplitude of the relevant seed projection (e.g., $f^{(1)}$ mapped to initial perturbation energy). After amplification the amplitude is approximately $G^{1/2}a_{0}$; transition to nonlinear cascade is plausible when $G^{1/2}a_{0}\gtrsim a_{\mathrm{NL}}$.
@@ 9 Remainder estimates and Chapman–Enskog error bounds
We state a model-precise remainder estimate.
@@
$C=C\big(T,\|(\rho,u,T)\|_{H^{s}_{x}},C_{\mathrm{inv}}\big),\qquad\varepsilon_{0}=\varepsilon_{0}\big(\|(\rho,u,T)\|_{H^{s}_{x}},C_{\mathrm{inv}}\big),$
@@ 10 Conclusion
Under explicit functional-analytic hypotheses we prove a deterministic operator-level mechanism by which first-order Chapman–Enskog kinetic corrections produce viscous constitutive behavior and enable macroscopic deviatoric stress in closed, unforced systems. The theorem asserts that production of $O(\varepsilon)$ deviatoric stress requires a nonzero $f^{(1)}$. A constructive BGK calculation reproduces the classical viscosity coefficient and demonstrates inversion and remainder control in an explicit setting. Extensions to boundary-driven or forced problems require boundary-layer analysis and are outlined for future work.
Acknowledgements
The author gratefully acknowledges Dr. Spyksma for his insightful and thorough review of the manuscript and Dr. Adam Barkman for his careful proofreading.
@@ Appendix A Pseudoinverse construction: full statement and proof sketch
$\langle-L_{x}g,g\rangle_{v,M}\geq\lambda_{0}\|g\|_{L^{2}_{v}(M^{-1})}^{2}.$
@@
We give a step-by-step constructive argument and indicate precise references where the black-box results are invoked.
Step 1: Frozen-$x$ operator and spectral gap. Fix $x_{0}$. Consider the frozen operator $L_{x_{0}}$ acting on $L^{2}_{v}(M(x_{0},\cdot)^{-1}dv)$. By hypothesis the spectral gap $\lambda_{0}>0$ on $N(L_{x_{0}})^{\perp}$ holds; therefore the restriction $L_{x_{0}}:N(L_{x_{0}})^{\perp}\to N(L_{x_{0}})^{\perp}$ is boundedly invertible. This is a standard consequence of coercivity; see Baranger & Mouhot [13, Thm. 1] or Mouhot & Strain [14] for constructive statements in kinetic settings.
@@
Remark on hypocoercivity: If the spectral gap hypothesis is replaced by a hypocoercivity framework (A5(b)), one obtains a bounded inverse in hypocoercive norms (with derivative losses) by following Villani’s constructive Lyapunov functional approach [15] and constructive implementations in [14]. The main difference is (i) $L^{-1}$ is bounded on a higher-regularity velocity space $H^{s_{v}}_{v}(M^{-1})$ and (ii) constants reflect derivative gains/losses required to close the hypocoercive estimates. ∎
@@ B.1 Equation for the remainder
$\partial_{t}R_{\varepsilon}+v\cdot\nabla_{x}R_{\varepsilon}-\frac{1}{\varepsilon}L_{x}R_{\varepsilon}=\mathcal{S}_{\varepsilon},$
@@ B.2 Projection and splitting
$R_{\varepsilon}=P_{x}R_{\varepsilon}+(I-P_{x})R_{\varepsilon}=:R_{\varepsilon}^{\parallel}+R_{\varepsilon}^{\perp}.$
@@
$\partial_{t}R_{\varepsilon}^{\perp}+v\cdot\nabla_{x}R_{\varepsilon}^{\perp}-\frac{1}{\varepsilon}L_{x}R_{\varepsilon}^{\perp}=(I-P_{x})\mathcal{S}_{\varepsilon}+\mathcal{C}_{\varepsilon},$
@@
$\|\mathcal{C}_{\varepsilon}\|\leq C_{\mathrm{comm}}\Big(\|R_{\varepsilon}^{\perp}\|+\|R_{\varepsilon}^{\parallel}\|\Big),$
@@ B.4 Moment equations for the parallel part
$\frac{d}{dt}R_{\varepsilon}^{\parallel}=\mathcal{M}R_{\varepsilon}^{\parallel}+\mathcal{F}_{\varepsilon},$
@@ B.5 Grönwall closure
$\frac{d}{dt}\|R_{\varepsilon}^{\perp}\|^{2}+\frac{2\lambda_{0}}{\varepsilon}\|R_{\varepsilon}^{\perp}\|^{2}\leq C_{5}\bigl(\|R_{\varepsilon}^{\perp}\|^{2}+\varepsilon^{4}+(\|R_{\varepsilon}^{\parallel}(0)\|+\varepsilon^{2})^{2}e^{2C_{4}t}\bigr).$
@@
$\|R_{\varepsilon}^{\perp}(t)\|^{2}\leq C_{6}\varepsilon^{4}+C_{7}\|R_{\varepsilon}^{\parallel}(0)\|^{2}e^{C_{8}t},$
@@
$\|R_{\varepsilon}(t)\|_{L^{2}_{x}L^{2}_{v}(M^{-1})}\leq C\varepsilon^{2},\qquad t\in[0,T],$

### v2

@@ 6 Main theorem
Here is the principal result in its operator-level form.
Theorem 6.1 (Molecular seed of shear: operator-level necessity).
Assume A1–A8 and the operator bounds of Lemmas 5.15.3.
Let $\{f_{\varepsilon}\}_{0<\varepsilon<\varepsilon_{0}}$ be sufficiently regular solutions of the rescaled Boltzmann equation under Chapman–Enskog scaling on the spatial domain specified in A2, and write the Chapman–Enskog expansion
@@
with $f^{(0)}(x,v,t)=M[\rho,u,T]$.
Assume the remainder satisfies the uniform bound
$\|R_{\varepsilon}\|_{L^{2}_{x}L^{2}_{v}(M^{-1})}\leq C\,\varepsilon^{2},$ (6.2)
for all $0<\varepsilon<\varepsilon_{0}$ in the norms and regularity framework of A3–A7.
All statements below are local in time on any interval $[0,T]$ on which the macroscopic fields remain in $H^{s}_{x}$ with the bounds required in A7.
Constants in (6.2)–(6.7) depend only on $\|(\rho,u,T)\|_{H^{s}_{x}}$, the spectral-gap or hypocoercivity constants, the uniform pseudoinverse bound $C_{\mathrm{inv}}$, and the domain geometry.
Uniqueness of $f^{(1)}$ below is always understood pointwise in $x$ in the weighted space of A3–A7, under the orthogonality condition (6.6).
Any initial-layer corrections violating (6.6) are absorbed into the parallel remainder and controlled by Section B.
Then:
@@
    (Constitutive decomposition).
    The pressure tensor admits the asymptotic expansion
@@
    The linear map $f^{(1)}\mapsto\tau^{(1)}$ is bounded under the velocity-moment weights of A3.
@@
    (Vanishing for uniform Maxwellians).
    If $f^{(0)}(x,v,t)$ is spatially uniform, then
    $(\partial_{t}^{(0)}+v\cdot\nabla_{x})f^{(0)}\equiv 0,$ (6.5)
    and under the Fredholm assumptions of A6 this implies $f^{(1)}(x,v,t)\equiv 0$ and thus $\tau^{(1)}\equiv 0$.
    Hence, in closed and unforced systems, no $O(\varepsilon)$ deviatoric stress can arise in the Chapman–Enskog expansion unless a nontrivial first correction is present.
@@
    (Uniqueness and explicit representation).
    Whenever the solvability orthogonality
@@
    holds for all collision invariants $\phi(v)$, the first correction is uniquely determined pointwise in $x$ by
@@
    where $L_{x}$ denotes the linearized collision operator at the local Maxwellian $M[\rho(x,t),u(x,t),T(x,t)]$.
    The mapping $f^{(0)}\mapsto f^{(1)}$ is continuous in the weighted norms of A3–A7, with operator norm bounded by $C_{\mathrm{inv}}$ from Lemma 5.2.
Corollary 6.2 (Deterministic necessity of the first kinetic correction).
@@ 7 BGK worked example
I illustrate the operator-level statements of Theorem 6.1 in the BGK setting.
Consider the BGK equation
@@
with constant relaxation time $\tau>0$.
The linearization at a local Maxwellian $M[\rho,u,T]$ gives
$L_{\mathrm{BGK}}g=-\frac{1}{\tau}\,(g-Pg),$ (7.2)
where $P$ denotes the orthogonal projection onto the nullspace $\mathrm{span}\{1,v_{i},|v|^{2}\}$.
Hence $L_{\mathrm{BGK}}^{-1}$ is explicit and satisfies
$\|L_{\mathrm{BGK}}^{-1}\|\leq\tau,$ (7.3)
so that $C_{\mathrm{inv}}=\tau$ in A6.
The Chapman–Enskog expansion is sought in the form
$f=M+\tau f^{(1)}+O(\tau^{2}),$ (7.4)
where the small parameter is $\varepsilon=\tau$.
At leading order $O(1)$, the equation yields $f^{(0)}=M$.
At order $O(\tau)$, the same first-order correction equation as in the Boltzmann case is obtained:
$(\partial_{t}^{(0)}+v\cdot\nabla_{x})M=-f^{(1)}.$ (7.5)
Thus the representation formula (5.6) reads simply
$f^{(1)}=-\,L_{\mathrm{BGK}}^{-1}\bigl((\partial_{t}^{(0)}+v\cdot\nabla_{x})M\bigr)=\tau\,(\partial_{t}^{(0)}+v\cdot\nabla_{x})M.$ (7.6)
Evaluation of the deviatoric stress
Fix a point $x_{0}$ and apply a Galilean shift so $u(x_{0})=0$.
The Maxwellian takes the standard form
@@
and I employ the classical Gaussian moment identities:
@@
The first-order stress is
$\tau^{(1)}_{ij}=\int_{\mathbb{R}^{d}}v_{i}v_{j}\,f^{(1)}(v)\,dv\\ =-\int_{\mathbb{R}^{d}}v_{i}v_{j}\,(\partial_{t}^{(0)}+v\cdot\nabla_{x})M(v)\,dv.$ (7.10)
Evaluating the derivatives of $M$ using the moment formulas above and writing all derivatives of $(\rho,u,T)$ at $x_{0}$ gives
$\tau^{(1)}_{ij}=-2\,\rho\,\tau\,R\,T\,S_{ij},$ (7.11)
where
$S_{ij}=\tfrac{1}{2}(\partial_{i}u_{j}+\partial_{j}u_{i})-\frac{1}{d}(\nabla\cdot u)\,\delta_{ij}$ (7.12)
is the deviatoric strain tensor.
Thus the standard BGK constitutive relation is recovered
@@
which matches the viscosity obtained from the exact BGK inversion $L^{-1}_{\mathrm{BGK}}$.
Remainder estimate
Since $L_{\mathrm{BGK}}^{-1}$ is explicit with operator norm $\tau$, the remainder estimate of A4 holds in the weighted space $L^{2}_{v}(M^{-1})$ with
$\|R_{\tau}\|_{L^{2}_{x}L^{2}_{v}(M^{-1})}\;\leq\;C(\|(\rho,u,T)\|_{H^{s}_{x}})\,\tau^{2}.$ (7.14)
The constant depends smoothly on $\|(\rho,u,T)\|_{H^{s}_{x}}$ and on the domain geometry but not on spatial position $x$, since the pseudoinverse is uniform.
Thus all hypotheses A1–A8 and Lemmas 5.15.3 are verified explicitly for the BGK model, and the operator-level necessity theorem applies directly.
@@ 8 Remainder estimates and Chapman–Enskog error bounds
I state a model-precise remainder estimate.
@@
$C=C\big(T,\|(\rho,u,T)\|_{H^{s}_{x}},C_{\mathrm{inv}}\big),\\ \qquad\varepsilon_{0}=\varepsilon_{0}\big(\|(\rho,u,T)\|_{H^{s}_{x}},C_{\mathrm{inv}}\big),$ (8.2)
@@ 9 Conclusion
Under explicit functional-analytic hypotheses I prove a deterministic operator-level mechanism by which first-order Chapman–Enskog kinetic corrections produce viscous constitutive behavior and enable macroscopic deviatoric stress in closed, unforced systems. The theorem asserts that production of $O(\varepsilon)$ deviatoric stress requires a nonzero $f^{(1)}$. A constructive BGK calculation reproduces the classical viscosity coefficient and demonstrates inversion and remainder control in an explicit setting. Extensions to boundary-driven or forced problems require boundary-layer analysis and are outlined for future work.
@@ Appendix A Pseudoinverse construction: full statement and proof sketch
$\langle-L_{x}g,g\rangle_{v,M}\geq\lambda_{0}\|g\|_{L^{2}_{v}(M^{-1})}^{2}.$ (A.1)
@@
I give a step-by-step constructive argument and indicate precise references where the black-box results are invoked.
Step 1: Frozen-$x$ operator and spectral gap. Fix $x_{0}$. Consider the frozen operator $L_{x_{0}}$ acting on $L^{2}_{v}(M(x_{0},\cdot)^{-1}dv)$. By hypothesis the spectral gap $\lambda_{0}>0$ on $N(L_{x_{0}})^{\perp}$ holds; therefore the restriction $L_{x_{0}}:N(L_{x_{0}})^{\perp}\to N(L_{x_{0}})^{\perp}$ is boundedly invertible. This is a standard consequence of coercivity; see Baranger & Mouhot [14, Thm. 1] or Mouhot & Strain [15] for constructive statements in kinetic settings.
@@
Remark on hypocoercivity: If the spectral gap hypothesis is replaced by a hypocoercivity framework (A5(b)), one obtains a bounded inverse in hypocoercive norms (with derivative losses) by following Villani’s constructive Lyapunov functional approach [16] and constructive implementations in [15]. The main difference is (i) $L^{-1}$ is bounded on a higher-regularity velocity space $H^{s_{v}}_{v}(M^{-1})$ and (ii) constants reflect derivative gains/losses required to close the hypocoercive estimates.
@@ B.1 Equation for the remainder
$\partial_{t}R_{\varepsilon}+v\cdot\nabla_{x}R_{\varepsilon}-\frac{1}{\varepsilon}L_{x}R_{\varepsilon}=\mathcal{S}_{\varepsilon},$ (B.1)
@@ B.2 Projection and splitting
$R_{\varepsilon}=P_{x}R_{\varepsilon}+(I-P_{x})R_{\varepsilon}=:R_{\varepsilon}^{\parallel}+R_{\varepsilon}^{\perp}.$ (B.2)
@@
$\partial_{t}R_{\varepsilon}^{\perp}+v\cdot\nabla_{x}R_{\varepsilon}^{\perp}-\frac{1}{\varepsilon}L_{x}R_{\varepsilon}^{\perp}=(I-P_{x})\mathcal{S}_{\varepsilon}+\mathcal{C}_{\varepsilon},$ (B.3)
@@
$\|\mathcal{C}_{\varepsilon}\|\leq C_{\mathrm{comm}}\Big(\|R_{\varepsilon}^{\perp}\|+\|R_{\varepsilon}^{\parallel}\|\Big),$ (B.4)
@@ B.4 Moment equations for the parallel part
$\frac{d}{dt}R_{\varepsilon}^{\parallel}=\mathcal{M}R_{\varepsilon}^{\parallel}+\mathcal{F}_{\varepsilon},$ (B.6)
@@ B.5 Grönwall closure
$\frac{d}{dt}\|R_{\varepsilon}^{\perp}\|^{2}+\frac{2\lambda_{0}}{\varepsilon}\|R_{\varepsilon}^{\perp}\|^{2}\\ \leq C_{5}\bigl(\|R_{\varepsilon}^{\perp}\|^{2}+\varepsilon^{4}+(\|R_{\varepsilon}^{\parallel}(0)\|+\varepsilon^{2})^{2}e^{2C_{4}t}\bigr).$ (B.8)
@@
$\|R_{\varepsilon}^{\perp}(t)\|^{2}\leq C_{6}\varepsilon^{4}+C_{7}\|R_{\varepsilon}^{\parallel}(0)\|^{2}e^{C_{8}t},$ (B.9)
@@
$\|R_{\varepsilon}(t)\|_{L^{2}_{x}L^{2}_{v}(M^{-1})}\leq C\varepsilon^{2},\qquad t\in[0,T],$ (B.10)